# Parity violation in atoms and electron scattering revisited

### v1

@@ Abstract
Exchange by two neutrinos (as well as by other fermions) generates a long-range parity-violating (PV) potential of the form $\sim G^{2}/r^{5}$, with characteristic range $\hbar/(2m_{\nu}c)$. This interaction produces a $0.8\%$ correction to the effective weak charge of the Cs atom, thereby resolving the $2\sigma$ discrepancy between the Standard Model prediction and the Cs PV measurement, and yielding $\sin^{2}\theta_{W}=0.2375(19)$ at $q^{2}\approx 0$ (to be compared with the Standard Model value $\sin^{2}\theta_{W}=0.2387$). We derive updated limits on an extra $Z^{\prime}$ boson and obtain a constraint on isospin-conserving oblique radiative corrections characterized by the Peskin–Takeuchi parameter, S=0.32(53)S=-0.32(53) at $q^{2}\approx 0$. The corresponding relative correction to the proton weak charge is $3\%$, and a similar correction is expected for the electron weak charge.
@@ I Introduction
Exchange by two neutrinos as in Fig. 1 produces long range potential $\sim G^{2}/r^{5}$ in the particle interaction, where $G$ is Fermi constant  [19, 18, 17, 25]. However, effects of parity conserving part of this potential are many orders of magnitude smaller than the sensitivity of the experiments [27, 1, 7, 46, 43, 41, 40, 32].
Figure 1: Feynman diagrams contributing to the long-range force due to neutrino exchange in atoms. Diagram (a) may also contain other fermions and quarks in the loop while the diagram (b) has only the electronic neutrino in the loop. Diagram (c) has been actually calculated at the quark level.
In Refs. [21, 24, 12, 23, 22] it was noted that the neutrino exchange potential has parity violating (PV) part $\sim G^{2}/r^{5}$ with the range $\hbar/2m_{\nu}c$ exceeding atomic size by several orders of magnitude. It was also demonstrated that mixed $Z-\gamma$ electron vacuum polarization produces PV potential $\sim\alpha G/r^{3}$ with the range $\hbar/2m_{e}c$=193 fm, which exceeds the range of the weak interaction by five orders of magnitude [21, 12], similar to the range of the Uehling potential due to electron vacuum polarization by the nuclear Coulomb field.
In the present paper we show that the contributions of the diagrams in Fig. 1 with exchange by two neutrinos or by other fermions to the PV effects in atoms is $\sim$1%, and this significantly exceeds the error 0.35% of the PNC measurement in Cs atom [49] and the error <<0.5% in the many-body atomic calculations of the $Z$-boson contribution [14, 10, 3, 4, 20, 34, 35, 9]. The work is in progress to improve both experimental [8, 44, 36] and theoretical [42] accuracy.
@@
Radiative corrections to the proton and neutron weak charges of the order $\alpha\approx 1/137$ have been calculated in Refs. [29, 28, 30, 5] and were shown to exceed 1%. However, diagrams on Fig. 1, which actually give comparable contributions, have not been included into these calculations. The point is that at large distance these diagrams give interaction proportional to $G^{2}$, which is very small. However, this interaction is highly singular, $\sim G^{2}/r^{5}$, and integration over distance with electron wave functions and effective cut-off at $r\sim 1/M_{Z}$ gives contribution $\sim G^{2}M_{Z}^{2}\sim\alpha G$. Thus, the suppression relative to the ordinary weak interaction mediated by the $Z$-boson exchange is $\sim\alpha$ [12].
@@ II Effective potential due to the fermion loop
where $W_{Z}$ and $W_{f}$ are the PV potentials produced by $Z$ boson tree-level exchenge and $f$ fermion loop correspondingly.
Dominating contribution to both matrix elements comes from the distance between electron and quark of order $r\sim 1/M_{Z}$, which is very small as compared to the scale of the variation of electron wave functions and nuclear size. Therefore, we may replace both potentials by their Fermi-type contact approximations $w_{Z,f}\delta(r)$ using relation $\int W_{Z,f}(r)4\pi r^{2}dr=w_{Z,f}\int\delta(r)d^{3}r$. Then the nuclear-spin-independent (NSI) part of the electron-nucleus PV operator $W(r)$ may be presented as
@@
where $Q_{W}=-N+(1-4\sin^{2}\theta_{W})Z$ is the weak nuclear charge, $\delta Q_{W}$ is the correction produced by the diagrams on Fig. 1, $\rho(r)$ is the nuclear density normalized by the condition $\int\rho(r)dV=1$, $N$ and $Z$ are numbers of neutrons and protons in the nucleus. Inclusion of radiative corrections from Refs. [29, 28, 30, 5] (see also [33, 16]) gives $Q_{W}=-0.98897N+(1.0250-3.9978\sin^{2}\theta_{W})Z$, which gives for the cesium atom the standard model value of the weak charge $Q_{W}^{SM}=-73.26(1)$, using $\sin^{2}\theta_{W}=0.23873$ for momentum $q^{2}\approx 0$ [16]. The experimental measurements of the parity violation in cesium atom give $\sin^{2}\theta_{W}=0.2349(18)$ [16].
@@
Calculations of the neutrino exchange potentials, presented in Appendix, are similar to that described in detail in Ref. [23]. Let us start from the diagram (a) in Fig. 1. The interaction constant on the bottom fermion line cancels with the single $Z$-boson exchange contribution in the ratio (1). Therefore, for any fermion on the bottom line (e.g. proton, neutron or electron in the electron- electron scattering case) the ratio of first diagram to the $Z$-boson contribution is the same. As a result we obtain
$\frac{\delta Q_{W}^{(1)}}{Q_{W}}=-\frac{g^{2}}{96\pi^{2}c_{W}^{2}}N_{\text{eff}}=-0.86\%\,$, (3)
where $g\approx 0.65$ is the $Z$-boson-neutrino interaction constant, $c_{W}^{2}=1-s^{2}_{W}=0.761$, $s_{W}^{2}\equiv\sin^{2}\theta_{W}$. Following Refs. [41, 12], we introduced here effective number of neutrinos $N_{\text{eff}}=14.5$. Indeed, there are 3 ordinary neutrinos. However, we should add to the loop all fermions with mass smaller than $Z$-boson mass, including $e,\,\mu,\,\tau,\,u,\,d,\,s,\,c,\,b$. $N_{\text{eff}}=14.5$ is not an integer number since interaction constants of other fermions with $Z$-boson are different from that for neutrinos.
Diagrams (b) and (c) in Fig. 1 do not have enhancement factor $N_{\text{eff}}$. Using the calculations of the momentum and radial integrals presented in Appendix with the result in Eq. (31), we obtain the following correction to the nuclear weak charge:
$\frac{\delta Q_{W}^{(2,3)}}{Q_{W}}=1.7\times 10^{-3}\frac{N-Z}{N-(1-4s^{2}_{W})Z}\,$, (4)
For the proton weak charge this ratio, enhanced by the small value of $(1-4\sin^{2}\theta_{W})=0.044$ in the denominator, is 3.8%. Adding contribution (3) corresponding to the diagram (a), we obtain the total correction to the weak charge
$\frac{\delta Q_{W}}{Q_{W}}=\left(-0.86+0.017\frac{N-Z}{N-(1-4s^{2}_{W})Z}\right)\%\,.$ (5)
Equation (5) shows that the contribution of the diagrams (b) and (c) (given in the last term) for all atoms, except for hydrogen, is small, so we obtain $\delta Q_{W}/Q_{W}\approx-0.8\%$ in all medium and heavy atoms. For the proton weak charge this gives,
@@ III Comparison with experimental data on weak charges and limits on new physics
If we use the value of the proton weak charge obtained in the recent very accurate lattice calculation, $Q_{W}^{SM}=0.06987(50)$, we obtain corrected value $Q_{W}^{SM}+\delta Q_{W}^{SM}=0.06987(1+0.03)=0.0719$. It agrees with $Q_{W}^{\text{exp}}=0.0719(45)$ measured by Qweak collaboration [26].
@@
The most sensitive point to the weak charge correction is currently in cesium PV experiments, where $\delta Q_{W}/Q_{W}\approx-0.8\%$ exceeds both the experimental error 0.35% and the error in our atomic many-body calculation ¡ 0.5%. This correction actually eliminates $2\sigma$ difference between the measured $Q_{W}^{\text{exp}}=-72.41(42)$ and the standard model $Q_{W}^{SM}+\delta Q_{W}^{SM}=-73.26(1-0.008)=-72.67$ weak charges.
@@
The difference $Q_{W}^{\text{exp}}-(Q_{W}^{SM}+\delta Q_{W}^{SM})=0.26(42)$ allows us to obtain limits on new physics. The limits on the additional $Z^{\prime}$-boson interaction strength obtained as a function of $Z^{\prime}$ mass from the measurements of PV in Cs atom are approximately 2 times stronger than that presented in Table II and Fig. 1 of Ref. [13].
@@ Appendix A Reduction of effective interaction potentials to the contact interaction
The exchange of various SM particles other than photon between the atomic nucleus and the electrons may be described by potential forces which may be considered as a perturbation to the Coulomb interaction potential. If such a potential $V(r)$ is short-range as compared with the size of the nucleus, they may be effectively reduced to the contact interaction with the the delta-function potential,
@@
where $g\approx 0.65$ is the gauge coupling constant of the $SU(2)_{L}$ group, and $c_{W}\equiv\cos\theta_{W}\approx 0.87$ is the cos of Weinberg angle. Since the range of this interaction is $r=O(m_{Z}^{-1})$, in atomic calculations it may be replaced with the contact interaction (9) with the coefficient
@@
Below, we consider effective potentials corresponding the exchange of neutrinos as per the Feynman graphs in Fig. 1. In the present work, we need two such potentials, which are usually referred to as the self-energy (SE) and ‘penguin’ (PG) contributions, respectively. These potentials were found in the works [22, 23]. We will calculate the corresponding contact interaction coefficients $\kappa$ for these potentials.
@@ A.1 Neutrino self-energy diagram
The potential corresponding to the neutrino one-loop self-energy diagram was calculated in Refs. [22, 23]:
@@
$\kappa_{\text{SE}}=4\pi\int_{0}^{\infty}r^{2}V(r)dr|_{m_{\nu}\to 0}=\frac{4\pi V_{0,\text{SE}}}{m_{Z}^{2}}\,.$ (18)
Here the following identity has been used:

### v2

# Effects of dispersion parity-violating interaction in electron scattering and atoms

@@ Abstract
Exchange of two neutrinos (as well as other fermions) generates a long-range parity-violating (PV) potential of the form $\sim G^{2}/r^{5}$, with characteristic range $\hbar/(2m_{\nu}c)$. In atomic systems the corresponding matrix elements converge at distances $r<10/M_{Z}$, so that the interaction between electron and quarks effectively reduces to a contact term $\sim G^{2}M_{Z}^{2}\delta^{3}(\vec{r})\sim G\,\alpha\,\delta^{3}(\vec{r})$. This interaction produces a $0.8\%$ correction to the effective weak charge of cesium, resolving the $2\sigma$ discrepancy between the Standard Model prediction and the measured Cs parity-violation amplitude. The corresponding value of the weak mixing angle is $\sin^{2}\theta_{W}=0.2375(19)$ at $q^{2}\approx 0$, in agreement with the Standard Model prediction $\sin^{2}\theta_{W}=0.23873$. The relative correction to the proton weak charge is about $3\%$, with a similar correction for the electron weak charge. Using these results, we revisit the limits on an additional $Z^{\prime}$ boson and obtain a constraint on isospin-conserving oblique radiative corrections characterized by the Peskin–Takeuchi parameter S=0.32(53)S=-0.32(53), at $q^{2}\approx 0$.
@@ I Introduction
Exchange of two neutrinos as in Fig. 1 produces long range potential $\propto G^{2}/r^{5}$ in the particle interaction, where $G$ is the Fermi constant [43, 20, 19, 18, 26]. The effects of parity conserving part of this potential are, however, many orders of magnitude smaller than the sensitivity of the experiments [28, 1, 7, 48, 45, 42, 41, 33] bf and are practically unobservable.
Figure 1: Feynman diagrams contributing to the long-range force due to neutrino exchange between electron $e$ and quark $q$.The self-energy (SE) diagram may also contain other leptons and quarks in the loop while the ‘penguin’ (PG1) diagram has only the electronic neutrino in the loop.
In Refs. [22, 25, 13, 24, 23] it was noted that the neutrino exchange potential has parity violating (PV) part $\propto G^{2}/r^{5}$ with the range $\hbar/2m_{\nu}c$ exceeding atomic size by several orders of magnitude. It was also demonstrated that mixed $\gamma Z$ electron vacuum polarization produces PV potential $\propto\alpha G/r^{3}$ with the range $\hbar/2m_{e}c=193$ fm, which exceeds the range of the weak interaction by five orders of magnitude [22, 13], similar to the range of the Uehling potential due to electron vacuum polarization by the nuclear Coulomb field.
In the present paper we show that the contribution of the diagrams in Fig. 1 with exchange by two neutrinos or by other fermions to the PV effects in atoms is $\sim$1%, and this significantly exceeds the error 0.35% of the PNC measurement in Cs atom [51] and the error <<0.5% in the many-body atomic calculations of the $Z$-boson contribution [15, 11, 3, 4, 21, 35, 36, 10]. The work is in progress to improve both experimental [8, 46, 37] and theoretical [44] accuracy.
@@
Radiative corrections to the proton and neutron weak charges of the order $\alpha\approx 1/137$ have been calculated in Refs. [30, 29, 31, 5, 9] and were shown to exceed 1%. However, diagrams on Fig. 1, which actually give comparable contributions, have not been included into these calculations. The point is that at large distance these diagrams give interaction proportional to $G^{2}$, which is very small. However, this interaction is highly singular, $\propto G^{2}/r^{5}$, and integration over distance with electron wave functions and effective cut-off at $r\sim 1/M_{Z}$ gives contribution $\sim G^{2}M_{Z}^{2}\sim\alpha G$. Thus, the suppression relative to the ordinary weak interaction mediated by the $Z$-boson exchange is $\sim\alpha$ [13].
@@
In this paper, we use natural units with $\hbar=c=1$.
@@ II Effective potential due to the fermion loop
where $W_{Z}$ and $W_{f}$ are the PV potentials produced by $Z$ boson tree-level exchange and $f$ fermion loop correspondingly.
Dominating contribution to both matrix elements comes from the distance between electron and quark of order $r\sim 1/M_{Z}$, which is very small as compared to the scale of the variation of electron wave functions and nuclear size. Therefore, we may replace both potentials by their Fermi-type contact approximations $w_{Z,f}\delta^{3}(\vec{r})$ using relation $\int W_{Z,f}(r)4\pi r^{2}dr=w_{Z,f}\int\delta^{3}(\vec{r})d^{3}r$. Then the nuclear-spin-independent (NSI) part of the electron-nucleus PV operator $W(r)$ may be cast in the form
@@
where $Q_{W}=-N+(1-4\sin^{2}\theta_{W})Z$ is the weak nuclear charge, $\delta Q_{W}$ is the correction produced by the diagrams on Fig. 1, $\rho(r)$ is the nuclear density normalized by the condition $\int\rho(r)d^{3}r=1$, $N$ and $Z$ are numbers of neutrons and protons in the nucleus, respectively. Inclusion of radiative corrections gives [34, 17]
$Q_{W}=$ $-0.98897N+(1.0250-3.9978\sin^{2}\theta_{W})Z$
$=$ $-0.98897N+0.070605Z\,.$ (3)
For cesium atom, this formula implies the standard model value of the weak charge $Q_{W}^{SM}=-73.26(1)$ assuming $\sin^{2}\theta_{W}=0.23873$ at zero momentum transfer [17]. The experimental measurements of the parity violation in cesium atom give $\sin^{2}\theta_{W}=0.2349(18)$ at $q=2.4$ MeV [17].
@@
Calculations of the neutrino exchange potentials, presented in Appendix, are similar to that described in detail in Ref. [24]. Let us start from the self-energy (SE) diagram in Fig. 1. The interaction constant on the bottom fermion line cancels with the single $Z$-boson exchange contribution in the ratio (1). Therefore, for any fermion on the bottom line (e.g. quark or electron in the electron-electron scattering case) the ratio of the SE diagram to the tree-level $Z$-boson contribution is the same. As a result we obtain
$\frac{\delta Q_{W}^{(\text{SE})}}{Q_{W}}=-\frac{g^{2}}{96\pi^{2}c_{W}^{2}}N_{\text{eff}}=-0.86\%\,$, (4)
where $g\approx 0.65$ is the $Z$-boson-neutrino interaction constant, $c_{W}^{2}=1-s^{2}_{W}=0.76127$, $s_{W}^{2}\equiv\sin^{2}\theta_{W}$. Following Refs. [42, 13], we introduced here effective number of neutrinos,
$N_{\text{eff}}\approx 14.5\,.$ (5)
Indeed, there are 3 ordinary neutrinos. However, we should add one-loop contributions of all fermions with masses smaller than that of the $Z$-boson, including $e,\,\mu,\,\tau,\,u,\,d,\,s,\,c,\,b$. $N_{\text{eff}}$ is not an integer number since interaction constants of other fermions with $Z$-boson are different from that for neutrinos, see Appendix B.1.
Diagrams PG1 and PG2 in Fig. 1 do not have enhancement factor $N_{\text{eff}}$. Using the calculations of the momentum and radial integrals presented in Appendix A.2 with the result in Eq. (34), we obtain the following correction to the nuclear weak charge (see also Appendix B.3 for details):
$\frac{\delta Q_{W}^{(\text{PG})}}{Q_{W}}=1.7\times 10^{-3}\frac{N-Z}{N-(1-4s^{2}_{W})Z}\,$, (6)
For the proton weak charge this ratio, enhanced by the small value of $(1-4\sin^{2}\theta_{W})=0.044$ in the denominator, is 3.8%. Adding contribution (4) corresponding to the SE diagram, we obtain the total correction to the nuclear weak charge
$\frac{\delta Q_{W}}{Q_{W}}=\left(-0.86+0.17\frac{N-Z}{N-(1-4s^{2}_{W})Z}\right)\%\,.$ (7)
Equation (7) shows that the contribution of the diagrams PG1 and PG2 (given in the last term) for all atoms, except for hydrogen, is small, so we obtain $\delta Q_{W}/Q_{W}\approx-0.8\%$ in all medium and heavy atoms. For the proton weak charge this gives,
@@
Distribution of proton and neutron densities in the nucleus are slightly different. Therefore to take into account this effect in numerical calculations it may be convenient to separate proton and neutron contributions to the weak charge. Adding the corrections (7) to Eq. (3) we find
$Q_{W}+\delta Q_{W}=$ $-0.98207N+(1.0181-3.9634s^{2}_{W})Z$
$=$ $-0.98207N+0.071918Z\,.$ (9)
Here we assume the SM value of the Weinberg angle with $\sin^{2}\theta_{W}=0.23873$ at zero momentum transfer [17].
@@ III Comparison with experimental data on weak charges and limits on new physics
If we use the value of the proton weak charge obtained in the recent very accurate lattice calculation [52], $Q_{W}^{SM}=0.06987(50)$, we obtain corrected value $Q_{W}^{SM}+\delta Q_{W}^{SM}=0.06987(1+0.03)=0.0719$. It agrees with $Q_{W}^{\text{exp}}=0.0719(45)$ measured by Qweak collaboration [27].
@@
The most sensitive point to the weak charge correction is currently in cesium PV experiments, where $\delta Q_{W}/Q_{W}\approx-0.8\%$ exceeds both the experimental error 0.35% and the error in our atomic many-body calculation which is under 0.5%. This correction actually eliminates $2\sigma$ difference between the measured $Q_{W}^{\text{exp}}=-72.41(42)$ and the standard model $Q_{W}^{SM}+\delta Q_{W}^{SM}=-73.26(1-0.008)=-72.67$ weak charges.
@@
The difference $Q_{W}^{\text{exp}}-(Q_{W}^{SM}+\delta Q_{W}^{SM})=0.26(42)$ allows us to obtain limits on new physics. The limits on the additional $Z^{\prime}$-boson interaction strength obtained as a function of $Z^{\prime}$ mass from the measurements of PV in Cs atom are approximately 2 times stronger than that presented in Table II and Fig. 1 of Ref. [14]. For $m_{Z^{\prime}}\gg\alpha Zm_{e}$, the products of $Z^{\prime}$ interaction constants with the electron and nucleus obeys the following bound: $|g_{e}^{A}g_{N}^{V}|/m_{Z^{\prime}}^{2}<2.2\times 10^{-8}$ GeV-2. For $m_{Z^{\prime}}\ll 1/R_{\text{atom}}$, the product of these constants is limited by $|g_{e}^{A}g_{N}^{V}|<1.7\times 10^{-14}$.
@@ Appendix A Reduction of effective interaction potentials to the contact interaction
The exchange of various SM particles other than photon between two point-like fermions may be described by potential forces which may be considered as a perturbation to the Coulomb interaction potential. If such a potential $V(r)$ is short-range as compared with the size of the nucleus, they may be effectively reduced to the contact interaction with the the delta-function potential,
@@
where $g\approx 0.65$ is the gauge coupling constant of the $SU(2)_{L}$ group, and $c_{W}\equiv\cos\theta_{W}\approx 0.87$ is the cos of Weinberg angle. We omit here Dirac matrix $\gamma_{5}$, see Eq. (2). Since the range of this interaction is $r=O(m_{Z}^{-1})$, in atomic calculations it may be replaced with the contact interaction (12) with the coefficient
@@
Below, we consider effective potentials corresponding to the exchange of neutrinos as per the Feynman graphs in Fig. 1. In the present work, we need two such potentials, which are usually referred to as the self-energy (SE) and ‘penguin’ (PG) contributions, respectively. These potentials were found in the works [23, 24]. We will calculate the corresponding contact interaction coefficients $\kappa$ for these potentials.
@@ A.1 Neutrino self-energy diagram
The potential corresponding to the massless neutrino one-loop self-energy diagram was calculated in Refs. [23, 24]:
@@
$\kappa_{\text{SE}}=4\pi\int_{0}^{\infty}r^{2}V(r)dr=\frac{4\pi V_{0,\text{SE}}}{m_{Z}^{2}}\,.$ (21)
Here the following identity was used:
@@ A.2 Triangle PG diagram contribution to the potential
Appendix B Calculation of vertex constant prefactor
In this paper, we study long-range forces between electron and nucleon with the range $r\gg O(m_{Z}^{-1})$. This interaction is mediated by SM particles with masses $m\ll m_{Z}$, which are $\nu^{\alpha}=(\nu_{e},\nu_{\mu},\nu_{\tau})$ for neutrinos, $l^{\alpha}=(e,\mu,\tau)$ for electron, muon and tau, $q_{k}=(u,d,s,c,b)$ for quarks. The relevant part of the SM Lagrangian is
${\cal L}=$ $\frac{g}{4c_{W}}[\bar{\nu}^{\alpha}\gamma^{\mu}(1-\gamma_{5})\nu^{\alpha}+\bar{l}^{\alpha}\gamma^{\mu}(c_{V}^{\alpha}-c_{A}^{\alpha}\gamma_{5})l^{\alpha}$
$+\bar{q}_{k}\gamma^{\mu}(c_{V}^{q_{k}}-c_{A}^{q_{k}}\gamma_{5})q_{k}]Z_{\mu}$
$+\frac{g}{2\sqrt{2}}[\bar{l}^{\alpha}\gamma^{\mu}(1-\gamma_{5})\nu^{\alpha}W^{-}_{\mu}$
$+\bar{u}\gamma^{\mu}(1-\gamma^{5})dW^{+}_{\mu}+h.c.]\,$ (35)
Here we ignore the the CKM matrix because the quarks in the loop may be effectively considered as massless.
The values of the vector and axial-vector constants are given in Table 1.
$c_{V}$ $c_{A}$
$l^{\alpha}$ $4s_{W}^{2}-1$ 1-1
$u,c,t$ $1-\frac{8}{3}s_{W}^{2}$ 1
$d,s,b$ $\frac{4}{3}s_{W}^{2}-1$ 1-1
$p$ $4s_{W}^{2} 11$
$n$ 1-1 1-1
Table 1: Values of the vector $c_{V}$ and axial-vector $c_{A}$ interaction constants for leptons $l^{\alpha}=(e,\mu,\tau)$, quarks, proton $p$ and neutron $n$. $s_{W}^{2}=\sin^{2}\theta_{W}$, where $\theta_{W}$ is the Weinberg angle.
B.1 Electron-proton effective interaction
At high momentum transfer, the proton is represented by three valence quarks $p=uud$. In the tree-level diagram, $Z$ boson interacts with each of these quarks, thus, $c_{V}^{p}=2c_{V}^{u}+c_{V}^{d}=1-4s_{W}^{2}$. Hence, the tree-level $Z$-boson exchange diagram contains the following combination of wave functions
$[\bar{e}\gamma^{\mu}(c_{V}^{e}-c_{A}^{e}\gamma_{5})e][\bar{p}\gamma_{\mu}(c_{V}^{\mu}-c_{A}^{\mu}\gamma_{5})p]$
$=-c_{A}^{e}c_{V}^{p}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{p}\gamma_{\mu}p]+\ldots\,$, (36)
and the corresponding prefactor is
$c_{A}^{e}c_{V}^{p}=1-4s_{W}^{2}\,.$ (37)
The same prefactor appears in the SE diagram in Fig 1. Therefore, these vertex constants cancel out in the ratio of matrix elements (1).
The SE diagram is, however, enhanced by the factor (5). In addition to the three neutrinos, this factor receives the following contributions. There are thee left leptons with the $Z$-boson interaction constant $\frac{1}{2}(c_{V}^{l}+c_{A}^{l})=2s_{W}^{2}-1$, and three right leptons with vertex constants $\frac{1}{2}(c_{V}^{l}-c_{A}^{l})=2s_{W}^{2}$. Then there are $2\times 3$ ($u$ and $c$ quarks with three colors) left up-like quarks with the $Z$-boson interaction constants $\frac{1}{2}(c_{V}^{u}+c_{A}^{u})=1-\frac{4}{3}s_{W}^{2}$, and the same number of the corresponding right quarks with the vertex constants $\frac{1}{2}(c_{V}^{u}-c_{A}^{u})=-\frac{4}{3}s_{W}^{2}$. Finally, there are $3\times 3$ ($d$, $s$, $b$ with three colors) right down-like quarks with the vertex constant $\frac{1}{2}(c_{V}^{d}+c_{A}^{d})=\frac{2}{3}s_{W}^{2}-1$, and the same number of the corresponding right quarks with the constant $\frac{1}{2}(c_{V}^{d}-c_{A}^{d})=\frac{2}{3}s_{W}^{2}$. All these coefficients enter the SE diagram in Fig. 1 quadratically. Combining them we find
$N_{\text{eff}}=\frac{160}{3}s_{W}^{4}-40s_{W}^{2}+21\approx 14.5\,.$ (38)
This explains the coefficients in Eq. (4).
The PG1 diagram in Fig. 1 has loop on the electron line, and $Z$-boson interacts with the proton line which consists of the three valence quarks. Therefore, this diagram contains the following combination of the wave functions (see also Appendix A in Ref. [24]):
$\bar{e}\gamma^{\mu}(1-\gamma_{5})e][\bar{p}\gamma_{\mu}(c_{V}^{p}-c_{A}^{p}\gamma_{5})p]=-c_{V}^{p}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{p}\gamma_{\mu}p]+\ldots$ (39)
There are three PG2 diagrams with the loop on the quark lines: two diagrams with $u\to d+W^{+}$, and one diagram with $d\to u+W^{-}$. These diagrams contain the following combinations of wave functions:
$[\bar{e}\gamma^{\mu}(-\gamma_{5}c_{A}^{e})e][\bar{d}\gamma_{\mu}c_{V}^{u}d]=$ $-c_{A}^{e}c_{V}^{u}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{d}\gamma_{\mu}d]$ (40)
$[\bar{e}\gamma^{\mu}(-\gamma_{5}c_{A}^{e})e][\bar{u}\gamma_{\mu}c_{V}^{d}u]=$ $-c_{A}^{e}c_{V}^{d}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{u}\gamma_{\mu}u]\,.$ (41)
The combined prefactor is
$c_{V}^{p}-2c_{A}^{e}c_{V}^{d}-c_{A}^{e}c_{V}^{u}=4s_{W}^{2}-1-1=-2+4s_{W}^{2}\,.$ (42)
The ratio of prefactors (42) and (37) is
$\frac{2-4s_{W}^{2}}{1-4s_{W}^{2}}\approx-23\,.$ (43)
The corresponding relative correction to the proton weak charge is the product of Eqs. (34) and (43):
$\frac{\delta Q_{p}}{Q_{p}}=-\frac{2-4s_{W}^{2}}{1-4s_{W}^{2}}\frac{\kappa_{\text{PG}}}{\kappa_{Z}}\approx\frac{1.7\times 10^{-3}}{1-4s_{W}^{2}}\,.$ (44)
B.2 Electron-neutron effective interaction
Neutron is represented by the three valence quarks, $n=udd$. The corresponding $Z$-boson vector coupling is $c_{V}^{n}=c_{V}^{u}+2c_{V}^{d}=-1$. Therefore, tree-level $Z$-boson exchange diagram has the prefactor
$c_{A}^{e}c_{V}^{n}=-1\,.$ (45)
The same prefactor appears in the SE diagram, which is enhanced by the effective number of neutrinos (5).
In the PG1 diagram, there is the following combination of wave functions:
$\bar{e}\gamma^{\mu}(1-\gamma_{5})e][\bar{n}\gamma_{\mu}(c_{V}^{n}-c_{A}^{n}\gamma_{5})n]=-c_{V}^{n}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{n}\gamma_{\mu}n]+\ldots$ (46)
There are also three PG2 diagrams: one diagram with $u\to d+W^{+}$, and two with $d\to u+W^{-}$. These diagrams contain the following combinations of wave functions:
$[\bar{e}\gamma^{\mu}(-\gamma_{5}c_{A}^{e})e][\bar{u}\gamma_{\mu}c_{V}^{d}u]=$ $-c_{A}^{e}c_{V}^{d}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{u}\gamma_{\mu}u]$ (47)
$[\bar{e}\gamma^{\mu}(-\gamma_{5}c_{A}^{e})e][\bar{d}\gamma_{\mu}c_{V}^{u}d]=$ $-c_{A}^{e}c_{V}^{u}[\bar{e}\gamma^{\mu}\gamma_{5}e][\bar{d}\gamma_{\mu}d].$ (48)
The combined prefactor is
$c_{V}^{n}-c_{A}^{e}c_{V}^{d}-2c_{A}^{e}c_{V}^{u}=2-4s_{W}^{2}\,.$ (49)
The ratio of prefactors (49) and (45) is
$\frac{2-4s_{W}^{2}}{-1}=-(2-4s_{W}^{2})\,.$ (50)
The corresponding relative correction to the neutron weak charge is the product of Eqs. (34) and (50):
$\frac{\delta Q_{n}}{Q_{n}}=-(2-4s_{W}^{2})\frac{\kappa_{\text{PG}}}{\kappa_{Z}}\approx 1.7\times 10^{-3}\,.$ (51)
B.3 Contribution to the nuclear weak charge due to PG diagrams
Combination of Eqs. (44) and (51) gives us the correction to the weak charge of a nucleus composed of $Z$ protons and $N$ neutrons due to the PG diagrams,
$\delta Q^{(\text{PG})}\approx\frac{1.7\times 10^{-3}}{1-4s_{W}^{2}}Z\cdot Q_{p}+1.7\times 10^{-3}N\cdot Q_{n}\,$, (52)
and substituting here tree-level weak charges $Q_{p}=1-4s_{W}^{2}$ and $Q_{n}=-1$, we have
$\delta Q^{(\text{PG})}\approx 1.7\times 10^{-3}(Z-N)\,.$ (53)
Dividing this expression by the tree-level value of the nuclear weak charge, $Q_{W}=-N+(1-4s_{W}^{2})Z$, we obtain the contribution (6).